OpenStudy (anonymous):

Find the zeros of the function in the interval [-2 pi, 2 pi] f(x) = 3cosx

OpenStudy (anonymous):

\[f(x)=0\iff 3\cos x=0 \\\iff\cos x=0 \\\iff x= \pm\frac{\pi}{2} \]